OpenStudy (anonymous):

As claudia walks away from a 264-cm lampost, the tip of her shadow moves twice as fast as she does. what is claudia's height?

OpenStudy (amistre64):

the key tends to be in the similar triangles

OpenStudy (amistre64):

is the tip of her shadow moving with respect to claude or to the lamppost?

OpenStudy (amistre64):

l/c = (d+s)/s l/c = d/s + 1 l/c - 1 = d/s s = d/(l/c-1) s = d/(l-c/c) s = cd/ (l-c) s' = cd'/(l-c) 2d' = cd'/(l-c) 2 = c/(l-c) 2(l-c) = c 2l-2c = c 2l = c+2c 2l = c(1+2) 2l/3 = c claud might be 2/3 the height of the lamppost